A rarefaction and an adjacent compression of a sound wave travelling in air are separated by a distance of 15cm. If the velocity of sound in Air is 330 m/s, What is the frequency?
1
Expert's answer
2019-01-15T09:04:11-0500

The frequency f of a sound wave is related to its velocity v by the equation f = v/λ, where λ is the wavelength. The wavelength is equal to the distance between neighboring rarefactions or between neighboring compressions, hence, it is twice the distance between a rarefaction and an adjacent compression. In our case then λ = 30 cm = 0.3 m, and f = (330/0.3) s–1 = 1100 s–1 = 1100 Hz.